# A Randomized Exchange Algorithm for Optimal Design of Multi-Response Experiments

### v1

@@ Abstract
Despite the increasing prevalence of vector observations, computation of optimal experimental design for multi-response models has received limited attention. To address this problem within the framework of approximate designs, we introduce mREX, an algorithm that generalizes the randomized exchange algorithm REX (J Am Stat Assoc 115:529, 2020), originally specialized for single-response models. The mREX algorithm incorporates several improvements: a novel method for computing efficient sparse initial designs, an extension to all differentiable Kiefer’s optimality criteria, and an efficient method for performing optimal exchanges of weights. For the most commonly used D-optimality criterion, we propose a technique for optimal weight exchanges based on the characteristic matrix polynomial. The mREX algorithm is applicable to linear, nonlinear, and generalized linear models, and scales well to large problems. It typically converges to optimal designs faster than available alternative methods, although it does not require advanced mathematical programming solvers. We demonstrate the application of mREX to bivariate dose-response Emax models for clinical trials, both without and with the inclusion of covariates.
@@
Optimal experimental design, Multi-response models, Convex optimization algorithms, D-optimality, Kiefer’s criteria, Emax models [^1] The research was supported by the Slovak Scientific Grant Agency (grant VEGA 1/0362/22) and the Collegium Talentum Programme of Hungary.
@@ 1 Introduction
where $\mathbf{H}(\mathbf{x}_{i})=\mathbf{f}(\mathbf{x}_{i})\mathbf{f}^{T}(\mathbf{x}_{i})$ is the “elementary” information matrix that measures the amount of information gained from one trial at the design point $\mathbf{x}_{i}\in\mathcal{X}$, for any $i$. For the standard model (2), the ranks of the elementary information matrices are $1$. The combined information matrix $\mathbf{M}(\mathbf{w})$ is clearly positive semidefinite, and can be of rank $m$.
@@
$\Phi_{p}\left[\mathbf{M}\right]=\begin{cases}\det\left[\mathbf{M}\right]^{1/m}&\text{ for }p=0\text{ and }\mathbf{M}\text{ nonsingular},\\ \left(\frac{1}{m}\>\mathrm{tr}\left[\mathbf{M}^{-p}\right]\right)^{-1/p}&\text{ for }p\in\left(0,\infty\right)\text{ and }\mathbf{M}\text{ nonsingular},\\ 0&\text{ for }\mathbf{M}\text{ singular}.\end{cases}$ (5)
@@
which is much less studied. Here $\mathbf{y}_{\mathbf{x}}$ represents the $s$-dimensional response from the trial at the design point $\mathbf{x}\in\mathcal{X}$, $\boldsymbol{\beta}\in\mathbb{R}^{m}$ is the vector of parameters, and $\mathbf{F}(\mathbf{x})$ is a known $m\times s$ matrix. The vectors $\boldsymbol{\epsilon}_{\mathbf{x}}$ of random errors are assumed to have expectation $\mathbf{0}_{s}$ and covariance matrix $\sigma^{2}\boldsymbol{\Sigma}$, where $\boldsymbol{\Sigma}$ is known and positive definite $s\times s$ matrix. Note that the dimension of $\boldsymbol{\Sigma}$ is typically small (we often consider bivariate response models, i.e., $s=2$), and it can be estimated from previous experiments (cf. [39], [1], [26]). Across different trials, the error vectors are assumed to be independent.
@@
Building on the work of [5], a randomized exchange algorithm (REX; [18]) was recently proposed for the single-response model. This algorithm rapidly converges even for problems with relatively large sizes $N$ of the design space and numbers $m$ of parameters, and outperforms the competing methods in many situations. In this paper, we propose the mREX (multi-response REX) algorithm, which generalizes REX to provide a fast and reliable solution for the multi-response settings.
@@
  • Extensions to involve all Kiefer’s $\Phi_{p}$-criteria with $p\in[0,\infty)$. Note that the original REX algorithm was formulated only for $D$- and $A$-optimality, corresponding to Kiefer’s criteria for $p=0$ and $p=1$, respectively. This extension allows experimenters to consider, for instance, criteria with $p\in(0,1)$ which form a compromise between $D$- and $A$-optimality, or criteria with larger $p$, which are close to the $E$-optimality criterion.
@@
We will show that the mREX algorithm retains the beneficial properties of the standard REX algorithm; in particular, it is simple to implement in any computing environment, can be applied to large problems, and generally outperforms the state-of-the-art competing methods applicable to the optimal designs problem of multi-response regression.
The rest of this paper is organized as follows. In Section 2, we summarize the fundamental mathematical properties of multi-response optimal design problem for the Kiefer’s optimality criteria, such as the lower bound on the efficiency of any given design. The central Section 3 presents the algorithm mREX, particularly the computation of initial designs and solution of the inner optimization problems. Section 4 examines extensions of the model (6) that allow for the use of mREX, and Section 5 provides numerical studies demonstrating the utility of mREX for an assessment of a dose-response design. In this section, we additionally demonstrate the computational advantages of mREX compared to existing algorithms capable of solving the multi-response optimal design problem. Proofs of mathematical theorems are given in Appendix.
@@ 2 The multi-response problem of Kiefer’s optimality
While the $\Phi_{p}$-optimal information matrix is unique, there is often a continuum of $\Phi_{p}$-optimal designs. Analogously to the single-response $\Phi_{p}$-optimal design problem, it is possible to show (cf. Section 8.2 in [31]) that there always exists a $\Phi_{p}$-optimal design with a support size of at most $\frac{1}{2}m(m+1)+1$. Hence, since in statistical applications we usually have $m\ll N$, there typically exists a “sparse” $\Phi_{p}$-optimal design. Importantly, while in the single-response case the size of the support of an optimal design is at least $m$, in the multi-response problem the optimal support size is often smaller than $m$. Clearly, a lower bound on the support size of any $\Phi_{p}$-optimal design is the minimum $k$ such that the matrix $(\mathbf{G}(\mathbf{x}_{i_{1}}),\ldots,\mathbf{G}(\mathbf{x}_{i_{k}}))$ has full row rank for some indices $i_{1},\ldots,i_{k}$, i.e., the lower bound is as small as $\lceil m/s\rceil$.
The $\Phi_{p}$-efficiency of a design $\mathbf{w}$ relative to a nonsingular design $\widetilde{\mathbf{w}}$ is $\Phi_{p}\left[\mathbf{M}(\mathbf{w})\right]/\Phi_{p}\left[\mathbf{M}(\widetilde{\mathbf{w}})\right]$. We refer to the $\Phi_{p}$-efficiency of $\mathbf{w}$ relative to the $\Phi_{p}$-optimal design $\mathbf{w}^{*}$ simply as $\Phi_{p}$-efficiency of $\mathbf{w}$ (e.g., Section 5.15 in [31]):
@@ 3 The randomized exchange algorithm for multi-response optimality
In this section, we detail our randomized exchange algorithm mREX; see Algorithm 1.
The original REX algorithm for a single-response model repeatedly updates the current design $\mathbf{w}$ by performing optimal (full or partial) exchanges of weights for pairs of points: one point $\mathbf{x}_{i^{s}}$ at which $\mathbf{w}$ is currently supported, and the other point $\mathbf{x}_{i^{g}}$ at which the so-called variance function is large, as such points are more promising to support an optimal design. In each iteration of the algorithm REX, the points $\mathbf{x}_{i^{s}}$ and $\mathbf{x}_{i^{g}}$ are selected in batches, and the sequence of pairings for optimal exchanges is randomized.
For our algorithm mREX, we generalize these steps to the multi-response settings, adding several new improvements. First, while the original REX is initialized with a uniformly random $m$-point design, for mREX we provide a more efficient initialization (Line 1 of Algorithm 1), the importance of which is even more pronounced for the multi-response optimal design problem, as there an optimal design is typically supported on fewer points than $m$. Second, we provide a method for selecting the design points for the exchanges (Line 4) that covers all Kiefer’s $\Phi_{p}$-criteria, not just $D$- and $A$-optimality as in REX. Third, since the optimal exchanges do not admit an explicit analytic solution in our case, we propose an efficient numerical method for computing optimal exchanges between candidate points (Line 8), particularly for the criterion of $D$-optimality.
Note that in the original REX algorithm, a tuning parameter $\gamma\geq 1/m$ controls the number of candidate points to exchange. More precisely, $L=\min(\lceil\gamma m\rceil,N)$ points are selected. We simplify the REX algorithm by setting $\gamma=1$ and additionally omitting the so-called leading Böhning exchange, which was required for the convergence proof of the REX algorithm. Our numerical experience suggests that even this simplified version converges rapidly to a $\Phi_{p}$-optimal design.
@@ 3.1 Initial design
to $\mathcal{S}$. In equation (14), $\mathbf{v}$ is a random direction in $\mathcal{C}^{\perp}(\mathbf{M}(\mathcal{S}))$, where $\mathbf{M}(\emptyset)=\mathbf{0}_{m\times m}$ and $\mathbf{M}(\mathcal{S})=\sum_{i\in\mathcal{S}}\mathbf{f}(\mathbf{x}_{i})\mathbf{f}^{T}(\mathbf{x}_{i})$ for $\mathcal{S}\neq\emptyset$. Geometrically, this corresponds to selecting a design point $\mathbf{x}_{i^{*}}$ such that $\mathbf{f}(\mathbf{x}_{i^{*}})$ is largest in a random direction in the subspace not yet spanned by $\{\mathbf{f}(\mathbf{x}_{i}):i\in\mathcal{S}\}$. The final output of the original KYM is a uniform design on $\{\mathbf{x}_{i}:i\in\mathcal{S}\}$; see [16] for details.
@@ 3.2 Candidate points
for $i=1,\ldots,N$, where $\Phi_{p}\left[\mathbf{M}(\mathbf{w})\right]$ and $\text{tr}\left[\mathbf{M}(\mathbf{w})^{-p}\right]$ are positive numbers, constant with respect to $i$. That is, the ordering of the design points given by $\partial\Phi_{p}\left[\mathbf{M}(\mathbf{w}),\mathbf{H}(\mathbf{x}_{i})\right]$, $i=1,\ldots,N$, is the same as the ordering given by the components of $\mathbf{g}_{p}(\mathbf{w})$, as defined in equation (12). This explains Line 4 of Algorithm 1. In fact, it is straightforward to verify that this generalizes the choice of the candidate support points in the original REX algorithm for the single-response problem and criteria of $D$- and $A$-optimality.
@@
for $i=1,\ldots,N$. Evaluating the sums of squares on the right-hand-side of (16) using available numerical matrix procedures can be much faster than computing the trace of $\mathbf{G}^{T}(\mathbf{x}_{i})\mathbf{M}^{-p-1}(\mathbf{w})\mathbf{G}(\mathbf{x}_{i})$, especially because it can be performed in a vectorized manner for all $i=1,\ldots,N$ (i.e., by avoiding explicit loops).
@@ 3.3 Optimal exchanges
For the single-response case, in which we always have $\mathrm{rank}(\mathbf{\Delta})\leq 2$, and for the criteria of $D$- and $A$-optimality, there exists an analytic solution of (17); see [5] for $D$-optimality and [18] for $A$-optimality. Recently, the analytic formulas for $D$-optimality have been extended to the case $\mathrm{rank}(\mathbf{\Delta})=3$; see [28]. Note that in Section 3.3.2 we show that for $D$-optimality a closed-form algebraic expression for $\alpha^{*}$ can in principle be derived if $\mathrm{rank}(\mathbf{\Delta})\leq 5$. However, the general multi-response case does not seem to lend itself to an analytical solution of (17), not even for the most common criteria. We therefore examine efficient numerical solutions.

### v2

@@ Abstract
Despite the increasing prevalence of vector observations, computation of optimal experimental design for multi-response models has received limited attention. To address this problem within the framework of approximate designs, we introduce mREX, an algorithm that generalizes the randomized exchange algorithm REX (J Am Stat Assoc 115:529, 2020), originally specialized for single-response models. The mREX algorithm incorporates several improvements: a novel method for computing efficient sparse initial designs, an extension to all differentiable Kiefer’s optimality criteria, and an efficient method for performing optimal exchanges of weights. For the most commonly used D-optimality criterion, we propose a technique for optimal weight exchanges based on the characteristic matrix polynomial. The mREX algorithm is applicable to linear, nonlinear, and generalized linear models, and scales well to large problems. It typically converges to optimal designs faster than available alternative methods, although it does not require advanced mathematical programming solvers. We demonstrate the usefulness of mREX on bivariate dose-response Emax models for clinical trials, both without and with the inclusion of covariates.
@@
Optimal experimental design, Multi-response models, Convex optimization algorithms, D-optimality, Kiefer’s criteria, Emax models [^2] The research was supported by the Collegium Talentum Programme of Hungary and the Slovak Scientific Grant Agency (grant VEGA 1/0327/25).
@@ 1 Introduction
where $\mathbf{H}(\mathbf{x}_{i})=\mathbf{f}(\mathbf{x}_{i})\mathbf{f}^{T}(\mathbf{x}_{i})$ is the “elementary” information matrix that measures the amount of information gained from one trial at the design point $\mathbf{x}_{i}\in\mathcal{X}$, for any $i$. For the standard model (2), the ranks of the elementary information matrices are $1$. The combined information matrix $\mathbf{M}(\mathbf{w})$ is clearly positive semidefinite, and can be of full rank $m$.
@@
$\Phi_{p}\left[\mathbf{M}\right]=\begin{cases}(\det\left[\mathbf{M}\right])^{1/m}&\text{ for }p=0\text{ and }\mathbf{M}\text{ nonsingular},\\ \left(\frac{1}{m}\>\mathrm{tr}\left[\mathbf{M}^{-p}\right]\right)^{-1/p}&\text{ for }p\in\left(0,\infty\right)\text{ and }\mathbf{M}\text{ nonsingular},\\ 0&\text{ for }\mathbf{M}\text{ singular}.\end{cases}$ (5)
@@
which is much less studied from the optimal design perspective. Here, $\mathbf{y}_{\mathbf{x}}$ represents the $s$-dimensional response from the trial at the design point $\mathbf{x}\in\mathcal{X}$, $\boldsymbol{\beta}\in\mathbb{R}^{m}$ is the vector of parameters, and $\mathbf{F}(\mathbf{x})$ is a known $m\times s$ matrix. The vectors $\boldsymbol{\epsilon}_{\mathbf{x}}$ of random errors are assumed to have expectation $\mathbf{0}_{s}$ and covariance matrix $\sigma^{2}\boldsymbol{\Sigma}$, where $\boldsymbol{\Sigma}$ is a known and positive definite $s\times s$ matrix. Note that the dimension of $\boldsymbol{\Sigma}$ is typically small (we often consider bivariate response models, i.e., $s=2$), and it can be estimated from previous experiments (cf. [39], [1], [26]). Across different trials, the error vectors are assumed to be independent.
@@
Building on the work of [5], a randomized exchange algorithm (REX; [18]) was recently proposed for the single-response model. This algorithm rapidly converges even for problems with relatively large sizes $N$ of the design space and numbers $m$ of parameters, and outperforms the competing methods in many situations (as demonstrated in Section 4 of [18]). In this paper, we propose the mREX (multi-response REX) algorithm, which generalizes REX to provide a fast and reliable solution for the multi-response settings.
@@
  • Extensions to involve all Kiefer’s $\Phi_{p}$-criteria with $p\in[0,\infty)$. The original REX algorithm was formulated only for $D$- and $A$-optimality, corresponding to Kiefer’s criteria for $p=0$ and $p=1$, respectively. This extension allows experimenters to consider, for instance, criteria with $p\in(0,1)$ which form a compromise between $D$- and $A$-optimality, or criteria with larger $p$, which are close to the $E$-optimality criterion.
@@
We will show that the mREX algorithm retains the beneficial properties of the standard REX algorithm; in particular, it is simple to implement in any computing environment, can be applied to large problems, and generally outperforms the state-of-the-art competing methods applicable to the optimal design problem of multi-response regression.
The rest of this paper is organized as follows. In Section 2, we summarize the fundamental mathematical properties of the multi-response optimal design problem for the Kiefer’s optimality criteria, such as the lower bound on the efficiency of any given design. The central Section 3 presents the algorithm mREX, particularly the computation of initial designs and solution of the inner optimization problems. Section 4 examines extensions of the model (6) that allow for the use of mREX, and Section 5 provides numerical studies demonstrating the utility of mREX for an assessment of a dose-response design. In this section, we additionally demonstrate the computational advantages of mREX compared to existing algorithms capable of solving the multi-response optimal design problem. Proofs of mathematical theorems are given in Appendix.
@@ 2 The multi-response problem of Kiefer’s optimality
While the $\Phi_{p}$-optimal information matrix is unique, there is often a continuum of $\Phi_{p}$-optimal designs. Analogously to the single-response $\Phi_{p}$-optimal design problem, it is possible to show (cf. Section 8.2 in [31]) that there always exists a $\Phi_{p}$-optimal design with a support size of at most $\frac{1}{2}m(m+1)+1$. Hence, since in statistical applications we usually have $m\ll N$, there typically exists a “sparse” $\Phi_{p}$-optimal design. Importantly, while in the single-response case the size of the support of an optimal design is at least $m$, in the multi-response problem the optimal support size is often smaller than $m$. Clearly, a lower bound on the support size of any $\Phi_{p}$-optimal design is the minimum $k$ such that the matrix $(\mathbf{G}(\mathbf{x}_{i_{1}}),\ldots,\mathbf{G}(\mathbf{x}_{i_{k}}))$ has full row rank for some indices $i_{1},\ldots,i_{k}$, i.e., the support of a $\Phi_{p}$-optimal design can be as small as $\lceil m/s\rceil$.
The $\Phi_{p}$-efficiency of a design $\mathbf{w}$ relative to a nonsingular design $\widetilde{\mathbf{w}}$ is $\Phi_{p}\left[\mathbf{M}(\mathbf{w})\right]/\Phi_{p}\left[\mathbf{M}(\widetilde{\mathbf{w}})\right]$. We refer to the $\Phi_{p}$-efficiency of $\mathbf{w}$ relative to a $\Phi_{p}$-optimal design $\mathbf{w}^{*}$ simply as $\Phi_{p}$-efficiency of $\mathbf{w}$ (e.g., Section 5.15 in [31]):
@@
Note that (11) aligns with its single-response counterpart used in the REX algorithm: In [18], the $D$- and $A$-efficiency of a nonsingular design $\mathbf{w}$ is bounded by
$\mathrm{eff}_{D}(\mathbf{w})\geq\frac{m}{\max_{i=1,\dots,N}(\mathbf{g}(\mathbf{w}))_{i}},$ (14)
$\mathrm{eff}_{A}(\mathbf{w})\geq\frac{\text{tr}\left(\mathbf{M}^{-1}(\mathbf{w})\right)}{\max_{i=1,\dots,N}(\mathbf{g}(\mathbf{w}))_{i}},$ (15)
where
$(\mathbf{g}(\mathbf{w}))_{i}=\begin{cases}\mathbf{f}^{T}(\mathbf{x}_{i})\mathbf{M}^{-1}(\mathbf{w})\mathbf{f}(\mathbf{x}_{i})\quad\text{ for $D$-optimality,}\\ \mathbf{f}^{T}(\mathbf{x}_{i})\mathbf{M}^{-2}(\mathbf{w})\mathbf{f}(\mathbf{x}_{i})\quad\text{ for $A$-optimality.}\end{cases}$ (16)
@@ 3 The randomized exchange algorithm for multi-response optimality
Algorithm 1 Randomized exchange algorithm for single-response models (REX) adapted from [18]: The value of $\mathrm{eff.bnd}(\mathbf{w})$ (Line 2) is the lower bound on the efficiency of the current design $\mathbf{w}$, as given by the right-hand side of (14) or (15). The vector $\mathbf{g}(\mathbf{w})$ (Line 4) is defined in equation (16) depending on whether we intend to compute $D$- or $A$-optimal designs.
[…]
2:Output: Approximate design $\mathbf{w}$
3:
4:Generate a random regular design $\mathbf{w}$
[…]
7:     Let $(i^{g}_{1},\ldots,i^{g}_{L})$ be a random permutation of the indices of the $L=\min(\lceil\gamma m\rceil,N)$ largest elements of $\mathbf{g}(\mathbf{w})$
8:     for $l\leftarrow 1$ to $L$ do
9:         for $k\leftarrow 1$ to $K$ do
10:              Set $\mathbf{\Delta}\leftarrow\mathbf{f}(\mathbf{x}_{i^{g}_{l}})\mathbf{f}^{T}(\mathbf{x}_{i^{g}_{l}})-\mathbf{f}(\mathbf{x}_{i^{s}_{k}})\mathbf{f}^{T}(\mathbf{x}_{i^{s}_{k}})$
11:              Find $\alpha^{*}$ in $\mathrm{argmax}\{\Phi[\mathbf{M}(\mathbf{w})+\alpha\mathbf{\Delta}]):\alpha\in[-w_{i^{g}_{l}},w_{i^{s}_{k}}]\}$
13:         end for
14:     end for
15:end while
@@
In this section, we detail our randomized exchange algorithm mREX. The original REX algorithm for a single-response model (Algorithm 1) repeatedly updates the current design $\mathbf{w}$ by performing optimal (full or partial) exchanges of weights for pairs of points: one point $\mathbf{x}_{i^{s}}$ at which $\mathbf{w}$ is currently supported, and the other point $\mathbf{x}_{i^{g}}$ at which the so-called variance function is large, as such points are more promising to support an optimal design. In each iteration of the algorithm REX, the points $\mathbf{x}_{i^{s}}$ and $\mathbf{x}_{i^{g}}$ are selected in batches, and the sequence of pairings for optimal exchanges is randomized.
For our algorithm mREX (Algorithm 2), we generalize these steps to the multi-response settings, adding several new improvements. First, the original REX is initialized with a uniformly random $m$-point design (Line 1 of Algorithm 1), possibly repeatedly, until the randomly generated design is nonsingular. For mREX, we provide a more efficient, nontrivial initialization (Line 1 of Algorithm 2, detailed in Section 3.1), the importance of which is even more pronounced for the multi-response optimal design problem, as there an optimal design is typically supported on fewer points than $m$. Second, we provide a method for selecting the design points for the exchanges (Line 4 of Algorithm 2) that covers all Kiefer’s $\Phi_{p}$-criteria, not just $D$- and $A$-optimality as in REX. Third, while the optimal exchanges in REX (Line 8 of Algorithm 1) are determined using analytic formulas, an explicit analytic solution is not available in the multi-response setting. Therefore, for mREX, we propose a computationally effective numerical approach (Line 8 of Algorithm 2, detailed in Section 3.3). Specifically, for the criterion of $D$-optimality, we reformulate the computation of optimal exchanges as finding the maximum of a fixed polynomial, which can be computed particularly efficiently. Note that the selection of candidate points, as well as the randomized-exchange approach itself, is generalized in a straightforward manner, while the approaches to computing the initial design and performing optimal exchanges between design points represent nontrivial extensions of the original REX algorithm.
In the original REX algorithm, a tuning parameter $\gamma\geq 1/m$ controls the number of candidate points to exchange. More precisely, $L=\min(\lceil\gamma m\rceil,N)$ points are selected. We simplify the REX algorithm by setting $\gamma=1$ and additionally omitting the so-called leading Böhning exchange, which was required for the convergence proof of the REX algorithm. Our numerical experience suggests that even this simplified version converges rapidly to a $\Phi_{p}$-optimal design.
@@ 3.1 Initial design
to $\mathcal{S}$. In equation (17), $\mathbf{v}$ is a random direction in $\mathcal{C}^{\perp}(\mathbf{M}(\mathcal{S}))$, where $\mathbf{M}(\emptyset)=\mathbf{0}_{m\times m}$ and $\mathbf{M}(\mathcal{S})=\sum_{i\in\mathcal{S}}\mathbf{f}(\mathbf{x}_{i})\mathbf{f}^{T}(\mathbf{x}_{i})$ for $\mathcal{S}\neq\emptyset$. Geometrically, this corresponds to selecting a design point $\mathbf{x}_{i^{*}}$ such that $\mathbf{f}(\mathbf{x}_{i^{*}})$ is largest in a random direction in the subspace orthogonal to the already selected $\{\mathbf{f}(\mathbf{x}_{i}):i\in\mathcal{S}\}$. The final output of the original KYM is a uniform design on $\{\mathbf{x}_{i}:i\in\mathcal{S}\}$; see [16] for details.
@@ 3.2 Candidate points
for $i=1,\ldots,N$, where $\Phi_{p}\left[\mathbf{M}(\mathbf{w})\right]$ and $\text{tr}\left[\mathbf{M}^{-p}(\mathbf{w})\right]$ are positive numbers, constant with respect to $i$. That is, the ordering of the design points given by $\partial\Phi_{p}\left[\mathbf{M}(\mathbf{w}),\mathbf{H}(\mathbf{x}_{i})\right]$, $i=1,\ldots,N$, is the same as the ordering given by the components of $\mathbf{g}_{p}(\mathbf{w})$, as defined in equation (12). This explains Line 4 of Algorithm 2. In fact, it is straightforward to verify that this generalizes the choice of the candidate support points in the original REX algorithm for the single-response problem and criteria of $D$- and $A$-optimality.
@@
for $i=1,\ldots,N$. Evaluating the sums of squares on the right-hand-side of (19) using available numerical matrix procedures can be much faster than computing the trace of $\mathbf{G}^{T}(\mathbf{x}_{i})\mathbf{M}^{-p-1}(\mathbf{w})\mathbf{G}(\mathbf{x}_{i})$, especially because it can be performed in a vectorized manner for all $i=1,\ldots,N$ simultaneously (i.e., by avoiding explicit loops).
@@ 3.3 Optimal exchanges
For the single-response case, in which we always have $\mathrm{rank}(\mathbf{\Delta})\leq 2$, and for the criteria of $D$- and $A$-optimality, there exists an analytic solution of (20); see [5] for $D$-optimality and [18] for $A$-optimality. Recently, the analytic formulas for $D$-optimality have been extended to the case $\mathrm{rank}(\mathbf{\Delta})=3$; see [28]. Note that in Section 3.3.2 we show that for $D$-optimality, a closed-form algebraic expression for $\alpha^{*}$ can in principle be derived if $\mathrm{rank}(\mathbf{\Delta})\leq 5$. However, the general multi-response case does not seem to lend itself to an analytical solution of (20), not even for the most common criteria. We therefore examine efficient numerical solutions.